OpenStudy (anonymous):

If tanA=b/a, where a and b are real numbers.Find sin^2A.

OpenStudy (anonymous):

@mathstudent55 - I got till that step! My doubt is that should i just leave it like that or can i simplifiy? Thanks.

OpenStudy (mathstudent55):

a^2 + b^2 = c^2 If tanA = b/a, then sinA = b/c sin^2A = b^2/c^2 sin^2A = b^2/(a^2 + b^2)

OpenStudy (mathstudent55):

I think this is it. The answer is in terms of a and b, like the given information.

OpenStudy (anonymous):

Thanks a lot!! :D

OpenStudy (mathstudent55):

wlcm